OpenStudy (anonymous):

Sketch the curve that satisfies the equation x^4 −4x^2 −x^2y^2 +4x^2 = 0 If you cannot give a sketch, see if you can find an argument as to why the solution curve is not the graph of a function.

OpenStudy (campbell_st):

collect like terms \[x^4 - x^2y^2 = 0\] then \[x^2y^2 = x^4\] divide both sides by x^2 \[y^2 = x^2\] take the square root of both sides \[y = \]

OpenStudy (campbell_st):

oops